# Diff-Shadow: Global-guided Diffusion Model for Shadow Removal

### v1

@@ 5 Conclusion
In this work, we propose a global-guided diffusion model, named Diff-Shadow, for high-quality shadow removal with no obvious boundaries and consistent illumination. We design a parallel network structure in the framework of the diffusion model to assist in the recovery of local patches using the shadow-free regions from the global image. In addition, we propose a sampling strategy named GSS to further improve the performance of shadow removal during diffusion denoising. Comprehensive experiments demonstrate the superiority of our Diff-shadow. which outperforms the SOTA methods on ISTD, ISTD+, and SRD datasets.
Appendix A Implementation Details
In all experiments, we employed a consistent network structure as described in Table 5. Our parallel network consists of two UNet structures.
@@ Appendix A Implementation Details
The local and global branches are nearly identical, with the only difference being the absence of a self-attention block in the global branch. The network incorporates six-scale resolutions and includes two residual blocks per-resolution. Detailed configurations and parameters can be found in Table 5.
@@
Table 5: Network-related configurations and parameter choices.
@@
Table 6: Diffusion-related configurations and parameter choices.
Appendix B More Visual Examples
Figure 7 presents some visual results on complex scenes to demonstrate the robustness of the proposed method. Figures 8910 illustrate some qualitative comparison with DHAN [4], DC-ShadowNet [20], Fu et al. [9], BMNet [49], and ShadowDiffusion [15] on ISTD+ [22] dataset. Our approach clearly outperforms previous methods by preserving the image texture and the luminance consistency. Figure 11 and Figure 12 exhibit some results of the proposed Diff-Shadow on SRD [31] and ISTD [41] datasets, respectively.
@@ Appendix B More Visual Examples
Figure 8: One example of shadow removal results on the ISTD+ [22] dataset. Comparison with DHAN [4], DC-ShadowNet [20], Fu et al. [9], BMNet [49] and ShadowDiffusion [15]. Corresponding values represent the ‘PSNR/SSIM’ scores. The best and second-best results are illustrated in red and blue.
Figure 9: One example of shadow removal results on the ISTD+ [22] dataset. Comparison with DHAN [4], DC-ShadowNet [20], Fu et al. [9], BMNet [49] and ShadowDiffusion [15] on ISTD+ [22] dataset. Corresponding values represent the ‘PSNR/SSIM’ scores. The best and second-best results are illustrated in red and blue.
Figure 10: One example of shadow removal results on the ISTD+ [22] dataset. Comparison with DHAN [4], DC-ShadowNet [20], Fu et al. [9], BMNet [49] and ShadowDiffusion [15]. Corresponding values represent the ‘PSNR/SSIM’ scores. The best and second-best results are illustrated in red and blue.
Figure 11: Shadow removal results of Diff-Shadow on the SRD [31] dataset.
Figure 12: Shadow removal results of Diff-Shadow on the ISTD [41] dataset.
Appendix C Limitations
The main limitation of our approach is the relatively long inference time compared to end-to-end image recovery networks and diffusion methods using simple UNet architectures. Our Diff-Shadow takes 8.72 seconds to recover an image of size $256\times 256$ on a single NVIDIA GTX 2080Ti GPU.
Another natural limitation of our model is the requirement for accurate shadow masks. This enables the network to efficiently use information from the non-shaded portion of the global image, and thus the accuracy of shadow detection directly determines the final result of the shadow removal. Similar to [14, 15, 18], we apply the masks from  [4] for fair comparisons, which is widely used on shadow removal tasks.

### v2

@@ Conclusion
In this work, we propose a global-guided diffusion model named Diff-Shadow, for high-quality shadow removal with no obvious boundaries and consistent illumination. We design a parallel network structure in the framework of the diffusion model to assist in the recovery of local patches using the shadow-free regions from the global image. In addition, we propose a sampling strategy named GSS to further improve the performance of shadow removal during diffusion denoising. Comprehensive experiments demonstrate that Diff-shadow outperforms the state-of-the-art.
Appendix A Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant Nos. 62372091, 62071097, 62402136 and in part by Sichuan Science and Technology Program under Grant Nos. 2023NSFSC0462, 2023NSFSC0458, 2023NSFSC1972.
Appendix B Diff-Shadow: Global-guided Diffusion Model for Shadow Removal Supplementary
In the supplementary material, we first provide more implementation details of the proposed Diff-Shadow.
Then, we present some visual results on complex scenes, and more visual comparisons on the ISTD+ (Le and Samaras 2019) and SRD (Qu et al. 2017) datasets to demonstrate the effectiveness of our approach. Further, we provide the analysis of differences with current methods. Finally, we discuss some limitations of the proposed method.
Appendix C Implementation Details
@@
In our implementation, we employed a consistent network structure, as described in Table 4. Our parallel network consists of two UNet structures. The local and global branches are nearly identical, with the only difference being the absence of a self-attention block in the global branch. The network incorporates six-scale resolutions and includes two residual blocks per-resolution. Detailed configurations and parameters can be found in Table 4.
@@
Appendix D More Results
Comparisons with ShadowDiffusion on Complex Dataset. Table 6 presents the quantitative comparison with ShadowDiffusion on the original resolution. In addition, the complex CUHK-Shadow dataset focuses on the segmentation of shadows and does not contain shadow-free ground truth. Figure 6 shows the visual results of generalizing the models trained on the SRD dataset to CUHK-Shadow dataset for testing. However, as shown in Figure 7, we also observed some failure cases. Noted, many shadow removal approaches struggle with such scenes because the shadows occupy relatively large areas of the image and are indistinguishable from the image content.
@@ Appendix D More Results
Table 4: Network configurations and parameter choices.
@@
Table 5: Diffusion configurations and parameter choices.
Resolution ShadowDiffusion Ours
$256\times 256$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
SRD 30.67 0.910 4.21 31.27 0.934 3.61
Table 6: RMSE values on SRD dataset.
Figure 6: Qualitative results on CUHK-Shadow dataset.
Figure 7: Failure case on CUHK-Shadow dataset.
[…]
(a) first-layer w/o 35.31
(b) w/o mid-layer 35.49
(c) second-layer mid-layer 36.27
(d) third-layer mid-layer 35.93
(e) first-layer&mid-layer w/o 35.40
(f) w/o every-layer 36.14
(h) first-layer mid-layer 36.47
Table 7: Ablation studies to evaluate the effectiveness of different network structures on the ISTD+ (Le and Samaras 2019) dataset.
Effectiveness of RCA module. We assess the impact of the RCA module and the Convolutional Aggregation (CA) at different layers on the ISTD+ (Le and Samaras 2019) dataset. Table 7 (a) and (b) demonstrate the lower performance when only the RCA module or only the CA module is applied. Table 7 (c) highlights the importance of the RCA module, showing its effectiveness in shallow feature fusion because low-level features capture essential elements, such as edges and colors, which are crucial for shadow removal tasks. Table 7 (d) confirms that adding the RCA module to both shallow and intermediate layers leads to unsatisfactory results due to limited mask size. Additionally, intermediate layer features primarily capture global semantic information, making them more suitable for fusion using the CA module. Table 7 (e) verifies the overall inferiority of solely incorporating the CA module at each layer compared to the full model in Table 7 (f) with the added RCA module. Therefore, the optimal approach is to apply the RCA module in the initial layer and the CA module in the intermediate layer.
@@
More Visual Results. Figure 8 presents some visual results on complex scenes to demonstrate the robustness of the proposed method. Figures 101112 illustrate some qualitative comparison with DHAN (Cun, Pun, and Shi 2020), DC-ShadowNet (Jin, Sharma, and Tan 2021), Fu et al. (Fu et al. 2021), BMNet (Zhu et al. 2022a), and ShadowDiffusion (Guo et al. 2023b) on ISTD+ (Le and Samaras 2019) dataset. Our approach clearly outperforms previous methods by preserving the image texture and the luminance consistency. Figure 13 and Figure 14 exhibit some results of the proposed Diff-Shadow on SRD (Qu et al. 2017) and ISTD (Wang, Li, and Yang 2018) datasets, respectively.
Figure 9: Qualitative comparison with original resolution for BMNet, ShadowFormer, and ours on ISTD (Le and Samaras 2019) dataset.
[…]
original size $640\times 480$ Input Image 20.33 0.874 11.35
DHAN (Cun, Pun, and Shi 2020) 27.88 0.921 6.29
BMNet (Zhu et al. 2022a) 29.02 0.928 4.17
ShadowFormer (Guo et al. 2023a) 30.47 0.935 4.79
Ours 30.58 0.938 3.69
Table 8: Quantitative results on images with original resolution from ISTD (Wang, Li, and Yang 2018) dataset.
Extension to Other Image Resolutions. Thanks to the patch-based diffusion modeling approach and the sampling process across overlapping patches during inference, the proposed method enables size-agnostic image restoration. As shown in Fig. 9, our Diff-Shadow maintains excellent illumination consistency even at the original image resolution ($640\times 480$) of the ISTD (Le and Samaras 2019) dataset. Furthermore, as illustrated in Table 8, the proposed Diff-Shadow also achieves state-of-the-art performances. Note that the comparison methods are trained on images of original resolution, while our training is based on images with a resolution of $256\times 256$.
@@
Figure 10: One example of shadow removal results on the ISTD+ (Le and Samaras 2019) dataset. Comparison with DHAN (Cun, Pun, and Shi 2020), DC-ShadowNet (Jin, Sharma, and Tan 2021), Fu et al. (Fu et al. 2021), BMNet (Zhu et al. 2022a) and ShadowDiffusion (Guo et al. 2023b). Corresponding values represent the ‘PSNR/SSIM’ scores. The best and second-best results are illustrated in red and blue.
Figure 11: One example of shadow removal results on the ISTD+ (Le and Samaras 2019) dataset. Comparison with DHAN (Cun, Pun, and Shi 2020), DC-ShadowNet (Jin, Sharma, and Tan 2021), Fu et al. (Fu et al. 2021), BMNet (Zhu et al. 2022a) and ShadowDiffusion (Guo et al. 2023b) on ISTD+ (Le and Samaras 2019) dataset. Corresponding values represent the ‘PSNR/SSIM’ scores. The best and second-best results are illustrated in red and blue.
Figure 12: One example of shadow removal results on the ISTD+ (Le and Samaras 2019) dataset. Comparison with DHAN (Cun, Pun, and Shi 2020), DC-ShadowNet (Jin, Sharma, and Tan 2021), Fu et al. (Fu et al. 2021), BMNet (Zhu et al. 2022a) and ShadowDiffusion (Guo et al. 2023b). Corresponding values represent the ‘PSNR/SSIM’ scores. The best and second-best results are illustrated in red and blue.
Figure 13: Shadow removal results of Diff-Shadow on the SRD (Qu et al. 2017) dataset.
Figure 14: Shadow removal results of Diff-Shadow on the ISTD (Wang, Li, and Yang 2018) dataset.
Appendix E Differences with Other Methods
Differences with WeatherDiffusion. The main differences include two aspects. First, in terms of network architecture, in contrast to WeatherDiffsuion’s (Özdenizci and Legenstein 2023) simple single-head UNet architecture, we adopt the parallel UNets with a local branch dedicated to noise estimation and a global branch designed for shadow-free image recovery. Additionally, we introduce the RCA module to avoid the problem of losing global information for patch recovery in WeatherDiffsuion. Second, regarding the sampling process, compared to the average weighting used in WeatherDiffsuion, our global guided sampling strategy (GSS) adaptively calculates patch fusion weights, ensuring the consistency of illumination in the shadow removal process. Finally, the ablation studies have demonstrated the considerable improvements.
Differences with CRFormer. First, in contrast to cross-attention modules in CRFormer (Wan et al. 2022b) whose $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ originates from the same global feature, the RCA integrates both local and global information. Furthermore, other attention-based methods boost the mutual information between shadow and non-shadow pairs through masks. However, the shadow information is dispensable for recovering non-shadow regions. Contrarily, our attention mechanism utilizes masks to suppress the global shadow values, integrating the non-shadow information from the global branch into the local branch. Finally, Table 9 shows the performance improvements compared with CRFormer.
CRFormer Ours
S NS ALL S NS ALL
ISTD+ 5.90 2.90 3.40 1.30 1.57 2.09
SRD 7.14 3.15 4.25 1.81 1.65 2.52
Table 9: RMSE values on ISTD+ and SRD datasets.
Appendix F Limitations
The main limitation of our approach is the relatively long inference time compared to end-to-end image recovery networks and diffusion methods using simple UNet architectures. Our Diff-Shadow takes 3.5 seconds (grid = 16) to recover an image of size $256\times 256$ on a single NVIDIA GTX 2080Ti GPU. The number of parameters in our method grows from 104M to 160M compared to the single-branch architecture, we believe that the increase in computation actually correlate well with improvements in results.
Another natural limitation of our model is the requirement for accurate shadow masks. This enables the network to efficiently use information from the non-shaded portion of the global image, and thus the accuracy of shadow detection directly determines the final result of the shadow removal. Similar to (Guo et al. 2023a, b; Hu et al. 2019), we apply the masks from  (Cun, Pun, and Shi 2020) for fair comparisons, which is widely used on shadow removal tasks.